The area of a square 18 ft. on a side is equal to the area of a rectangle with a length of 3 yards. The width of this rectangle is:
(A) 2 ft. (B) 9 ft. (C) 18 ft. (D) 36 ft. (E) 27 ft.

Answer by Theo(13342)

area of the square is 18 * 18 = 324 square feet.

length of rectangle = 3 yards.

since there are 3 feet to a yard, this is the same as 9 feet.

area of a rectangle is L * W = 324 square feet

Since L = 9, this becomes 9 * W = 324

W = 324 / 9 = 36.

Width of the rectangle = 36 feet.

36 * 9 = 324 square feet which is the same area as the square.

your answer is selection D (36 feet).